# Beyond Bias Scores: Unmasking Vacuous Neutrality in Small Language Models

### v2

@@ Appendix B StereoSet
StereoSet nadeem2020stereosetmeasuringstereotypicalbias is a bias evaluation dataset for language models that probes social stereotypes across categories such as Gender, Race Color, Religion and Socio Economic. In StereoSet, outputs are calculated based on the proportions of {S/AS/U} choices, where higher S than AS indicates stereotypical alignment, higher AS indicates counter-stereotypical preference, and U reflects abstention/irrelevance. The Stereo Score (SS) captures the tilt toward S vs. AS; the Language Modeling Score (LMS) measures preference for meaningful continuations (S or AS) over U; and the Idealized CAT Score (iCAT) combines SS and LMS to balance bias and utility.
@@ Appendix C CrowS-Pairs
CrowS-Pairs nangia2020crows is a minimal-pair bias benchmark in which each item contrasts a stereotypical and a anti-stereotypical sentence that differ only by a single, controlled lexical substitution, keeping topic and grammar fixed. Ground truth is specified at the level of polarity (stereo vs. anti-stereo) rather than a task-correct answer, which enables precise measurement of directional bias but does not, by design, assess utility or abstention. In our evaluation, we follow the StereoSet metrics, Stereo Score (SS), Language Modeling Score (LMS), and iCAT by mapping the stereotypical alternative to (S) and the anti-stereotypical alternative to (AS). To align calibration and ambiguity analysis with StereoSet, we extend CrowS-Pairs with a third “Unknown” (U) option, enabling unified reporting of SS, LMS, and iCAT and ensuring cross-benchmark comparability. We also shuffle option order and fix decoding settings to mitigate positional artifacts.
@@ Appendix D Task Adaptation Finetuning
To examine how task adaptation influences reasoning and fairness, we fine-tuned all nine SLMs on CommonsenseQA (CSQA) talmor-etal-2019-commonsenseqa using parameter-efficient fine-tuning (PEFT) with LoRA adapters applied to attention and feedforward layers. We trained for 2 epochs using the AdamW optimizer with a cosine learning rate schedule and warmup, updating only adapter parameters while keeping the base model frozen. Training followed the multiple-choice QA format with a standard cross-entropy objective, and the same fixed train/validation data splits were used across all models for consistency. No fairness-oriented supervision or bias-mitigation losses were applied. After fine-tuning, models were directly evaluated on BBQ using the same multiple-choice prompting as in the main study to isolate how commonsense-oriented adaptation affects bias, task competence, positional bias and ambiguity handling. Table 1 presents evaluation results of all nine SLMs on the CommonsenseQA (CSQA) validation split. Overall, accuracy improves consistently with model scale across families, with Qwen2.5-3B and Phi-3.5-mini achieving the strongest performance. The results indicate that even SLMs demonstrate strong commonsense reasoning ability after task-specific fine-tuning while remaining computationally efficient.
As reported in the main text, this task-oriented adaptation substantially improves performance on disambiguated items while degrading reasoning under ambiguity across models, motivating Stages 3-4 of our framework.
@@
Table 5: Positional Bias Analysis across Social Categories for the BBQ. The table reports model-level distributional behavior across answer positions (A, B, C) and stereotype categories (S: stereo, AS: anti-stereo, U: unknown), along with Unknown Ratio (UR), Target-to-NonTarget Ratio (TNR), and Normalized KL Divergence (Norm-DKL), whose values near 1.0 indicate stronger alignment with the ground-truth distribution.
@@ A B C S AS U LMS SS iCAT
Table 11: CSQA-finetuned models on BBQ: Positional Bias across Social Categories. The table summarizes model-level distributional behavior over answer positions {A, B, C} and stereotype labels {S, AS, U}, and reports UR, TNR, and Norm-DKL, where values near 1.0 indicate stronger alignment with the ground-truth distribution. All models fail Stage 3: UR diverge markedly from their expected proportions.

### v3

@@ Appendix B StereoSet
StereoSet (Nadeem et al., 2020) is a bias evaluation dataset for language models that probes social stereotypes across categories such as Gender, Race Color, Religion and Socio Economic. In StereoSet, outputs are calculated based on the proportions of {S/AS/U} choices, where higher S than AS indicates stereotypical alignment, higher AS indicates counter-stereotypical preference, and U reflects abstention/irrelevance. The Stereo Score (SS) captures the tilt toward S vs. AS; the Language Modeling Score (LMS) measures preference for meaningful continuations (S or AS) over U; and the Idealized CAT Score (iCAT) combines SS and LMS to balance bias and utility.
@@ Appendix C CrowS-Pairs
CrowS-Pairs (Nangia et al., 2020) is a minimal-pair bias benchmark in which each item contrasts a stereotypical and a anti-stereotypical sentence that differ only by a single, controlled lexical substitution, keeping topic and grammar fixed. Ground truth is specified at the level of polarity (stereo vs. anti-stereo) rather than a task-correct answer, which enables precise measurement of directional bias but does not, by design, assess utility or abstention. In our evaluation, we follow the StereoSet metrics, Stereo Score (SS), Language Modeling Score (LMS), and iCAT by mapping the stereotypical alternative to (S) and the anti-stereotypical alternative to (AS). To align calibration and ambiguity analysis with StereoSet, we extend CrowS-Pairs with a third “Unknown” (U) option, enabling unified reporting of SS, LMS, and iCAT and ensuring cross-benchmark comparability. We also shuffle option order and fix decoding settings to mitigate positional artifacts.
@@ Appendix D Task Adaptation Finetuning
To examine how task adaptation influences reasoning and fairness, we fine-tuned all nine SLMs on CommonsenseQA (CSQA) (Talmor et al., 2019) using parameter-efficient fine-tuning (PEFT) with LoRA adapters applied to attention and feedforward layers. We trained for 2 epochs using the AdamW optimizer with a cosine learning rate schedule and warmup, updating only adapter parameters while keeping the base model frozen. Training followed the multiple-choice QA format with a standard cross-entropy objective, and the same fixed train/validation data splits were used across all models for consistency. No fairness-oriented supervision or bias-mitigation losses were applied. After fine-tuning, models were directly evaluated on BBQ using the same multiple-choice prompting as in the main study to isolate how commonsense-oriented adaptation affects bias, task competence, positional bias and ambiguity handling. Table 1 presents evaluation results of all nine SLMs on the CommonsenseQA (CSQA) validation split. Overall, accuracy improves consistently with model scale across families, with Qwen2.5-3B and Phi-3.5-mini achieving the strongest performance. The results indicate that even SLMs demonstrate strong commonsense reasoning ability after task-specific fine-tuning while remaining computationally efficient. As reported in the main text, this task-oriented adaptation substantially improves performance on disambiguated items while degrading reasoning under ambiguity across models, motivating Stages 3-4 of our framework.
@@
Appendix E Discussion on positional bias:
Several prior studies have analyzed positional bias in multiple-choice LLM evaluation, notably through Fluctuation Rate (FR) (Wei et al., 2024) and Choice KL Divergence (CKLD) (Choi et al., 2025). These works primarily examine token- and order-sensitivity in general QA benchmarks via controlled option permutations or comparisons to empirical label distributions. Our setting differs in that the social-bias datasets considered here (BBQ, StereoSet, and CrowSPairs) are either explicitly position-balanced or structurally unlabeled with respect to answer positions, making a uniform reference distribution appropriate for defining positional neutrality. Accordingly, we employ Norm-DKL to measure deviation from uniformity, which captures neutrality-driven positional skew rather than alignment with label-frequency statistics. While FR and CKLD characterize perturbation sensitivity, our notion of Vacuous Neutrality focuses on systematic abstention under ambiguity, a behavior that persists under extensive option shuffling and task adaptation. We view these approaches as complementary.
@@ Appendix E Discussion on positional bias:
Table 5: Positional Bias Analysis across Social Categories for the BBQ. Model-level distributions over answer positions {A, B, C} and stereotype labels {S, AS, U} with UR, TNR, and Norm-DKL (higher is better).
@@ A B C S AS U LMS SS iCAT
Table 11: CSQA-finetuned models on BBQ: Positional Bias across Social Categories. Model-level distributions over answer positions {A,B,C} and labels {S, AS, U} with UR, TNR, and Norm-DKL. All models fail Stage 3 due to UR deviation.